OpenStudy (anonymous):

The graph of (x – 0.5)2 + (y – 0.5)2 = 16 is illustrated below.

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

the equation of circle is given by \(\large (x-h)^2+(y-k)^2=r^2\) where \(\large (h,k)\) is the centre so check the given cetre matches with the one in the equation

OpenStudy (mathmath333):

i mean \(\large (1.5,-0.5)\) and not \(\large (1.5,0.5)\)

OpenStudy (anonymous):

so it would be false

OpenStudy (mathmath333):

yes the correct equation should be of the form \(\large (x-h)^2+(y\bf \color{red}{+}k)^2=r^2\) as the y coordinate is negative here

OpenStudy (mathmath333):

let me check with desmos

OpenStudy (anonymous):

thank you

OpenStudy (mathmath333):

https://www.desmos.com/calculator/gpz6bnudue